OpenStudy (anonymous):

The length of a rectangle is 3m less than twice its width, and the area of the rectangle is 27m^2 . Find the dimensions of the rectangle.

OpenStudy (anonymous):

i already solved once !!

OpenStudy (anonymous):

it said it was wrong

OpenStudy (campbell_st):

let w = width thne length = 2w - 3 the area is 27 = w(2w -3) so you are solving a quadratic \[2w^2 - 3w - 27 = 0\] just solve for w

OpenStudy (anonymous):

w=3,9/2

OpenStudy (anonymous):

whats the length

OpenStudy (campbell_st):

well you are almost there the factored form is (2w - 9)(w + 3) = 0 measurements can't be zero so only consider the positive solution.

OpenStudy (campbell_st):

oops measurements can't be negative.... only consider the positive solution.